Please help me solve this equation. The question is find two consecutive odd integers such that the sum of their reciprocals is 12/35.
Answer by scott8148(6628)

(1 / n) + [1 / (n + 2)] = 12 / 35

35n + 70 + 35n = 12n^2 + 24n

0 = 12n^2 - 46n - 70

0 = 6n^2 - 23n - 35

0 = (6n + 7) (n - 5)

n = 5 ___ n + 2 = 7